Two objects each of mass 1.5 kg, are moving in the same straight line but in opposite directions. The velocity of each object is 2.5 ms–1 before the collision during which they stick together. What will be the velocity of the combined object after collision?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Mass of the objects m1 = m2 = 1.5 kg
Velocity of first object v1= 2.5 m/s
Velocity of second object v2= – 2.5 m/s
Momentum before collision = m1v1= m2v2
= (1.5 × 2.5) + (1.5 × –2.5)
= 0
Momentum after collision
= m1+ m2= 1.5 + 1.5
= 3.0 kg
After collision v = ?
According to law of conservation of momentum
Momentum before collision = Momentum after collision
0 = 3 × v
v = 0


Thanks